# references.bib

### v1

@@ Abstract
Uncooperative satellite engagements with nation-state actors prompts the need for enhanced maneuverability and agility on-orbit. However, robust, autonomous and rapid adversary avoidance capabilities for the space environment is seldom studied. Further, the capability constrained nature of many space vehicles does not afford robust space situational awareness capabilities that can inform maneuvers. We present a “Cat & Mouse” system for training optimal adversary avoidance algorithms using Reinforcement Learning (RL). We propose the novel approach of utilizing intercepted radio frequency communication and dynamic spacecraft state as multi-modal input that could inform paths for a mouse to outmaneuver the cat satellite. Given the current ubiquitous use of RF communications, our proposed system can be applicable to a diverse array of satellites. In addition to providing a comprehensive framework for an RL architecture capable of training performant and adaptive adversary avoidance policies, we also explore several optimization based methods for adversarial avoidance on real-world data obtained from the Space Surveillance Network (SSN) to analyze the benefits and limitations of different avoidance methods.
@@
Simulation, Multi-Agent, Reinforcement Learning, RF, Communications, Localization, Space, Adversary
@@ I Introduction
In March of 2025, Chinese satellites exhibited dog-fighting capabilities [McCarthy_2025], following years of both Russian [Hitchens_2023] and Chinese [Jones_2023a] satellites approaching dangerously close to US satellites in geosynchronous orbit. Such uncooperative activity prompts the need for satellite agility and maneuverability which can be facilitated through edge-based autonomy. To achieve this, appropriate sensing would be required to properly characterize the contested environment. Not all satellites have precise space domain awareness (SDA) sensing suites onboard, despite having powerful buses and flight controllers that can facilitate autonomous operations. We propose leveraging an uncooperative space vehicle’s communication systems as a means to evaluate safe flight control policies to safely navigate contested domains in situations where support from the ground is not feasible.
Contested environment scenarios when engaging with an intelligent adversary in space is poorly studied in the open literature. In part, this is a function of the historically poor maneuverability and sensing in space which is necessary for intelligent evasion. Recent investments in the ‘new space’ industry have made such maneuverability more commonplace, thereby calling for novel intelligence capabilities to provide autonomous decision insight for these space vehicles. We propose a feedback architecture for uncooperative maneuvering that involves coupling an array of communications transceivers and a centralized reinforcement learning model to train a spacecraft to tactically maneuver around a non-cooperative satellite. We call our proposed scenario a ”Cat & Mouse” system which entails a “mouse” spacecraft that attempts to evade a second “cat” satellite’s sensing capabilities while the “cat” pursues the “mouse”. This generalized scenario is informed by recent events[Hitchens_2023, Jones_2023a, McCarthy_2025]. The intention is to use the cat’s communications capabilities sensed in lower altitude orbits to estimate the cat’s location and train a complex behavioral policy for contested space environments using a Reinforcement Learning (RL) approach as illustrated in Figure 1.
@@
Our motivation for utilizing RF sensors is to develop a system that engages sensors ubiquitous across satellites. This is in contrast to other sensing modalities such as optical or laser-ranging hardware which is only found on a subset of satellites. Moreover, the practice of locating unknown RF sources is well-studied, and several methods can be engaged to equip a nominally designed satellite with necessary data to autonomously maneuver in contested environments. Using these insights, we can more specifically define our Cat & Mouse scenario. We assume that the cat and mouse are sufficiently far apart such that the mouse cannot sense the cat with on-board sensors. Additionally, we assume that the mouse spacecraft can be aided by receiving additional information from satellites in lower orbit intercepting the cat’s RF communication signals. The mouse must then use this information to compute optimal spacecraft control.
@@
  • Train a Reinforcement Learning policy which incorporates both the dynamic spacecraft state as well as the processed RF observations in order to provide optimal evasive control for the mouse spacecraft
@@ II Related Works
In recent years there has been growing attention focused on contested space, and satellite behavior when confronted with adversarial spacecraft. [prince2018, doi:10.2514/1.A35800, yang2024]. However, these works often use optimization and heuristic based solutions which are only applicable within the narrow range of their target scenario, and lack the ability to develop display behaviors. Furthermore, autonomous capabilities in space are severely bottelnecked by our current space infrastructure’s Space Sensing Awareness (SSA) capabilities, which cannot necessarily provide exact position or orbit of the adversary in real time [8569343]. If autonomous satellite capabilities are to improve, new methods for sensing and detecting non-cooperative satellites must be explored.
@@
In addition to potential sensing modalities for SSA, it is important to recognize the current standards for satellite planning and their limitations. The majority of proposed works frame satellite maneuvering as a complex optimization problems that assumes accurate telemetry data of the satellite as well as its environment [8187728]. Luo et. al. proposes a novel adaptive differential evolutionary algorithm for optimizing earth-observation satellite maneuvers [rs14091966]. Porcelli et. al. provide a method for detecting and estimating satellite maneuvers using radar observations assuming an observation accuracy of 5 meters [PORCELLI2022274]. Hu et al. proposes a complex mathematical method for path planning in space environments with close proximity to obstacles, but again require full knowledge of the satellites surrounding state [Huxie]. While all of the discussed proposed methods satisfy the requirements for specific space missions, their computational complexity and inability to handle noisy observations highlights their inability to perform on the edge and with RF informed observations which can be significantly noisier than radar, laser, or optics.
@@
Our work proposes a new model for edge based satellite autonomy which addresses shortcomings both in SSA and modern satellite maneuvering algorithms by exploring new sensing modalities and the ability of RL policies to provide robust and optimal control. This method leverages RF transmissions intercepted in LEO to estimate the location of a potential adversary. We then show that this observational data can be utilized by a Reinforcement Learning policy to provide highly reactive and optimized control, keeping the potential adversary at a desired distance, while minimizing fuel cost and deviation from the initial orbit.
III Simulation
In order to train a successful RL policy a simulation environment that accurately emulates the contested space environment as well as the sensor data collected is essential. In this section we define the framework for our simulation which models the orbital dynamics of each spacecraft, sensor constellation configurations, and finally RF localization.
III-A Orbital Dynamics Modeling
The dynamic modeling of our simulation is composed of two layers: first the orbital dynamics of all satellites can be modeled using Kepler’s Laws of Planetary Motion, second the localized dynamics of the cat and mouse spacecraft are modeled using the Clohessy Wiltshire Equations. In order to do this the dynamics of the system are described in several reference frames: first, the inertial frame or Earth-Centered, Earth-Fixed (ECEF) notated by $\hat{\text{{I}}},\hat{\text{{J}}},\hat{\text{{K}}}$, second, the orbital frame notated by $\hat{\text{{x}}}_{o},\hat{\text{{y}}}_{o},\hat{\text{{z}}}_{o}$, and finally, the localized frame or Hill frame notated by $\hat{\text{{x}}}_{H},\hat{\text{{y}}}_{H},\hat{\text{{z}}}_{H}$. An image depicting the three frames and their relations can be seen in Figure 2. In simulation, coordinates are generally kept in the ECEF and Hill frame, translating between the two frames can be done by:
@@ III-A Orbital Dynamics Modeling
All orbits are assumed to be near circular, and therefore each one can be defined by five parameters: inclination ($i$), argument of periapsis ($\omega$), right ascension of accending node (RAAN) ($\Omega$), semi-major axis ($a$), and mean anomaly (M). Additionally, it should be noted that in Equations 2 and 3 the notation $\boldsymbol{c}_{\Omega}$ represents $cos(\Omega)$.
@@
Equations 4-9 allow us to solve for the position and velocity of the a spacecraft’s orbit at any point in time during the simulation.
In a localized frame the state of our cat and mouse spacecraft can be described using the Clohessy Wiltshire (CW) Equations. These equations model the behavior of a spacecraft in relation to a single point moving on a near circular orbit. We can track the location of this point using Equations 4-9. The equations of motion for the two spacecrafts position relative to this point is then:
@@
III-B RF Modeling and Constellation Configuration
@@
$\Delta d_{i}=||\boldsymbol{x}^{E}_{a}-\boldsymbol{x}^{E}_{i}||-||\boldsymbol{x}^{E}_{a}-\boldsymbol{x}^{E}_{0}||$ (15)
In Equation 15 $\boldsymbol{x}^{E}_{0}$ is the location of the reference sensor.
III-C Cat Localization
@@
Multiple works cited in Section II have shown their methods to perform similar to the CRLB both in simulation and real world testing [ZHOU2021109758, s23146254, 7582404, 10356280]. Thus, in order to reduce computational complexity of the simulation the CRLB is computed using Equations 16-19 in order to provide an appropriate noise level for the cat’s location estimate rather than run a TDOA localization algorithm using simulated TDOA data every timestep. The corresponding position estimate can be simulated by summing a sampled error using the CRLB and the true position of the cat:
${\boldsymbol{x}^{E}_{a}}^{\prime}=\boldsymbol{x}^{E}_{a}+\mathcal{N}(0,\,\sqrt{diag(\text{CRLB})})$ (20)
Due to the distance between the RF sensors and source as well as the fact that the sensors are constantly moving in relation to the source the size of the noise can vary from nominal to extreme. A graph depicting the error of the position estimates during a single episode can be seen in Figure 4. For this reason, we find it necessary to incorporate a filter into our system in order to properly process the observation data which will be discussed in Section IV.
@@
IV Method
In this section we will discuss our RL training and implementation architecture. In order to properly evade an uncooperative actor while maintaining minimal deviation from the initial path an agent must be capable of perceiving its surroundings, processing observations, computing optimal actions, and finally executing said actions in a single feedback loop. In the system proposed noisy raw observation data is received from the TDOA localization algorithm. This data is then filtered using an Extended Kalman Filter (EKF) to reduce noise, and the fed into an RL model to provide optimal control in the form of positional commands. Finally, the necessary thrusts are computed by an Model Predictive Controller (MPC) in order to be applied in simulation. We will discuss each step of this in greater detail bellow.
IV-A Extended Kalman Filter
As mentioned, the position estimates provided by the TDOA localization algorithm can be extremely noisy. In order to mitigate this, an EKF is implemented to smooth the position data before it is fed into the RL model. The EKF implemented follows a architecture often implemented in space systems [springerSatelliteOrbits]. At each timestep, the received position estimates are updated by estimating the expected location due to the dynamics of the system and previous observation, as well as observed noise of the system:
${\boldsymbol{x}^{E}_{a,i}}^{+}={\boldsymbol{x}^{E}_{a,i}}^{-}+\boldsymbol{K}_{i}({\boldsymbol{x}^{E}_{a,i}}^{\prime}-{\boldsymbol{x}^{E}_{a,i}}^{-})$ (21)
where ${\boldsymbol{x}^{E}_{a}}^{+}$ is the updated state, ${\boldsymbol{x}^{E}_{a}}^{-}$ is the predicted current state given the previous updated state, and $\boldsymbol{K}_{i}$ is the Kalman gain. In order to predict the future state we utilize Cowell’s Formulation, which propagates the previous known state of the system the using Newtons Law of Universal Gravitation:
$\boldsymbol{F}=\frac{Gm_{1}m_{2}}{r^{3}}\hat{\boldsymbol{r}}$ (22)
The updated state of the cat ${\boldsymbol{x}^{E}_{a,i}}^{+}$ is then used as an observation of the current state of the system for the RL model. A more in depth description of the EKF implemented can be found in the appendix.
IV-B Reinforcement Learning
Once the observational data is filltered, it is used as input to an RL policy tasked with computing optimal positional control. The observation provided to the cat is a concatenation of the mouse’s state in the local frame, the mouse’s current action, and the estimated state of the cat in the local frame from the last $N$ timesteps:
$\boldsymbol{o}_{i}=[\boldsymbol{x}^{H}_{m,i},\boldsymbol{v}^{H}_{m,i},a_{i-1},{\boldsymbol{x}^{H}_{a,i-N}}^{+},...,{\boldsymbol{x}^{H}_{a,i}}^{+}]$ (23)
It should be noted that we found the EKF provides the greatest benefit when implemented during testing in combination with an RL policy that was trained on un-filtered data. Thus, during training the observation space is:
$\boldsymbol{o}_{i}=[\boldsymbol{x}^{H}_{m,i},\boldsymbol{v}^{H}_{m,i},a_{i-1},{\boldsymbol{x}^{H}_{a,i-N}},...,{\boldsymbol{x}^{H}_{a,i}}]$ (24)
The resulting action $\boldsymbol{a}_{i}$ is the desired change in current location of the mouse. This is then used to compute the desired location of the mouse in the local coordinate frame, and an MPC is utilized to compute the necessary thrust control to be applied. During training, the RL policy is evaluated by a reward function which prioritizes deviation from the initial orbit, fuel consumed, and proximity to the cat:

### v2

@@ Abstract
This work presents a novel method for leveraging intercepted Radio Frequency (RF) signals to inform a constrained Reinforcement Learning (RL) policy for robust control of a satellite operating in contested environments. Uncooperative satellite engagements with nation-state actors prompts the need for enhanced maneuverability and agility on-orbit. However, robust, autonomous and rapid adversary avoidance capabilities for the space environment is seldom studied. Further, the capability constrained nature of many space vehicles does not afford robust space situational awareness capabilities that can be used for well informed maneuvering. We present a “Cat & Mouse” system for training optimal adversary avoidance algorithms using RL. We propose the novel approach of utilizing intercepted radio frequency communication and dynamic spacecraft state as multi-modal input that could inform paths for a mouse to outmaneuver the cat satellite. Given the current ubiquitous use of RF communications, our proposed system can be applicable to a diverse array of satellites. In addition to providing a comprehensive framework for training and implementing a constrained RL policy capable of providing control for robust adversary avoidance, we also explore several optimization based methods for adversarial avoidance. These methods were then tested on real-world data obtained from the Space Surveillance Network (SSN) to analyze the benefits and limitations of different avoidance methods.
@@
On-Orbit Simulation, Multi-Agent, Reinforcement Learning, constrained Reinforcement Learning, RF, Communications, Localization, Space, Adversary
@@ I Introduction
In March of 2025, Chinese satellites exhibited dog-fighting capabilities [McCarthy_2025], following years of both Russian [Hitchens_2023] and Chinese [Jones_2023a] satellites approaching dangerously close to US satellites in geosynchronous orbit. Such uncooperative activity prompts the need for satellite agility and maneuverability which can be facilitated through edge-based autonomy. To achieve this, appropriate sensing would be required to properly characterize the contested environment. Not all satellites have precise space domain awareness (SDA) sensing suites onboard, despite having powerful buses and flight controllers that can facilitate autonomous operations. We propose leveraging an uncooperative space vehicle’s communication systems as a means to evaluate safe flight control policies to carefully navigate contested domains in situations where support from the ground is not feasible.
Contested environment scenarios when engaging with an intelligent adversary in space is seldom discussed in the open literature. In part, this is a function of the historically poor maneuverability and sensing in space which is necessary for intelligent evasion. Recent investments in the ‘new space’ industry have made such maneuverability more commonplace, thereby calling for novel intelligence capabilities to provide autonomous decision insight for these space vehicles. We propose a feedback architecture for uncooperative maneuvering that involves coupling an array of communications transceivers and a centralized reinforcement learning model to train a spacecraft to tactically maneuver around a non-cooperative satellite. We call our proposed scenario a ”Cat & Mouse” system which entails a “mouse” spacecraft that attempts to evade a second “cat” satellite’s sensing capabilities while the “cat” pursues the “mouse”. This generalized scenario is informed by recent events[Hitchens_2023, Jones_2023a, McCarthy_2025]. The intention is to use the cat’s communications capabilities sensed by collaborative assets in lower altitude orbits to estimate the cat’s location and train a complex behavioral policy for contested space environments using a Reinforcement Learning (RL) approach as illustrated in Figure 1.
@@
Our motivation for utilizing RF sensors is to develop a system that engages a ubiquitous sensing modality across all satellites. This is in contrast to other traditional space situational awareness sensors such as optical or laser-ranging hardware which is only found on a subset of satellites. Moreover, the practice of locating unknown RF sources is well-studied, and several methods can be engaged to equip a nominally designed satellite with necessary data to autonomously maneuver in contested environments. Using these insights, we can more specifically define our cat & mouse scenario. We assume that the cat and mouse are sufficiently far apart such that the mouse cannot sense the cat with on-board sensors. Additionally, we assume that the mouse spacecraft can be aided by receiving additional information from satellites in lower orbit intercepting the cat’s RF communication signals. The mouse must then use this information to compute optimal spacecraft control.
@@
  • Design a Constrained Reinforcement Learning policy which incorporates both the dynamic spacecraft state as well as the processed RF observations in order to provide optimal evasive control for the mouse spacecraft
@@ II Related Works
In recent years there has been growing attention focused on contested space, and satellite behavior when confronted with adversarial spacecraft. [prince2018, doi:10.2514/1.A35800, yang2024]. However, these works often use optimization and heuristic based solutions which are only applicable within the narrow range of their target scenario, and lack the ability to develop display behaviors. Furthermore, autonomous capabilities in space are severely bottelnecked by our current space infrastructure’s Space Sensing Awareness (SDA) capabilities, which cannot necessarily provide exact position or orbit of the adversary in real time [8569343]. If autonomous satellite capabilities are to improve, new methods for sensing and detecting non-cooperative satellites must be explored.
@@
In addition to potential sensing modalities for space domain awareness (SDA), it is important to recognize the current standards for satellite planning and their limitations. The majority of proposed works frame satellite maneuvering as a complex optimization problems that assumes accurate telemetry data of the satellite as well as its environment [8187728]. Luo et. al. proposes a novel adaptive differential evolutionary algorithm for optimizing earth-observation satellite maneuvers [rs14091966]. Porcelli et. al. provide a method for detecting and estimating satellite maneuvers using radar observations assuming an observation accuracy of 5 meters [PORCELLI2022274]. Hu et al. proposes a complex mathematical method for path planning in space environments with close proximity to obstacles, but again require full knowledge of the satellites surrounding state [Huxie]. While all of the discussed proposed methods satisfy the requirements for specific space missions, their computational complexity and inability to handle noisy observations highlights their inability to perform on the edge and with RF informed observations which can be significantly noisier than radar, laser, or optics.
@@
Our work proposes a new model for edge based satellite autonomy which addresses shortcomings both in SDA and modern satellite maneuvering algorithms by exploring new sensing modalities and the ability of RL policies to provide robust and optimal control. This method leverages RF transmissions intercepted in LEO to estimate the location of a potential adversary. We then show that this observational data can be utilized by a Reinforcement Learning policy to provide highly reactive and optimized control, keeping the potential adversary at a desired distance, while minimizing fuel cost and deviation from the initial orbit.
III Preliminaries and Problem Definition
We will first describe the preliminary information required to frame the RF informed evasive planning problem. This first entails a formal definition of the multiple frames necessary to describe the behavior and motion of the multiple types of spacecraft involved in the problem statement. Then, we provide a brief overview of Markov Decision Processes (MDPs) and the framing of the proposed problem statement as such.
III-A Reference Frames
In order to model the problem, the dynamics of the system are described in several reference frames: first, the inertial frame or Earth-Centered, Earth-Fixed (ECEF) notated by $\hat{\text{{I}}},\hat{\text{{J}}},\hat{\text{{K}}}$, second, the orbital frame notated by $\hat{\text{{x}}}_{o},\hat{\text{{y}}}_{o},\hat{\text{{z}}}_{o}$, and finally, the localized frame or Hill frame notated by $\hat{\text{{x}}}_{H},\hat{\text{{y}}}_{H},\hat{\text{{z}}}_{H}$. An image depicting the three frames and their relations can be seen in Figure 2. In simulation, coordinates are generally kept in the ECEF and Hill frame, translating between the two frames can be done by:
@@ III-A Reference Frames
where $i,\omega,\Omega$ are the orbital angles inclination, argument of periapsis, and right ascension of ascending node (RAAN) respectively. And, $M$ is the mean anomaly of the orbit. Additionally, it should be noted that in Equations 2 and 3 the notation $\boldsymbol{c}_{\Omega}$ represents $cos(\Omega)$. Each spacecraft can be tracked in all three frames at once, however, spacecraft that are assumed not to have thrust control such as LEO constellation assets are only described in the ECEF and orbital frame.
@@
III-B MDP Problem Definition
In order to solve for the sequential decision making process required for reactive evasive maneuvering we model the system as an infinite horizon MDP which provides a framework for describing stochastic environments with nondeterministic actions by state-action pairs [bellman1, bellman2]. An infinite horizon MDP can be formally defined by $\mathcal{M}=(\mathcal{S},\mathcal{A},T,s_{0},R,\gamma)$, where $\mathcal{S}$ is the state space of the environment, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ is the transition function, $s_{0}:\mathcal{S}$ is the initial state, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ is the reward function, and $\gamma\in[0,1)$ is a discount factor. At each discrete timestep, an agent receives the current state of the environment $s$, and follows a policy $\pi_{\phi}$ to chose an action $a$. The agent then receives a reward $r$ and is informed of the resulting state $s^{\prime}$. An optimal policy for the MDP $\pi_{\phi}^{*}:\mathcal{S}\times\mathcal{A}$ chooses actions that maximize the expected discounted utility:
$\pi_{\phi}^{*}=\underset{a\in\mathcal{A}}{\operatorname{argmax}}Q^{*}(s,a)$ (4)
where $Q^{*}(s,a)$ is the expected utility of taking action $a$ at state $s$. Critical to learning the optimal policy is mapping the transition function $T(s,a,s^{\prime})\forall s\in\mathcal{S},a\in\mathcal{A},s^{\prime}\in\mathcal{S}$, which provides the probability of transitioning to state $s^{\prime}$ given you take action $a$ from state $s$. The dynamics of both the cat and mouse spacecraft are modeled as linear and time discrete:
$\begin{bmatrix}\boldsymbol{x}^{H}_{i+1}\\ \dot{\boldsymbol{x}}^{H}_{i+1}\end{bmatrix}=A\begin{bmatrix}\boldsymbol{x}^{H}_{i}\\ \dot{\boldsymbol{x}}^{H}_{i}\end{bmatrix}+B\begin{bmatrix}\boldsymbol{0}\\ \boldsymbol{u}_{i}\end{bmatrix}$ (5)
where $\boldsymbol{x}\in\mathbb{R}^{3}$ represents the position of the spacecraft in cartesian coordinates, $\dot{\boldsymbol{x}}\in\mathbb{R}^{3}$ represents the translational velocities, and $\boldsymbol{u}_{i}\in\mathbb{R}^{3}$ is the control computed by $f(\boldsymbol{x}_{i},\dot{\boldsymbol{x}}_{i},\boldsymbol{a}_{i}):\mathbb{R}^{9}\rightarrow\in\mathbb{R}^{3}$ which converts a desired change in position $\boldsymbol{a}_{i}$ to thrust control $\boldsymbol{u}_{i}$. Therefore, the full state at any point can be described by a timeseries of recent cat position estimates, dynamic state of the mouse spacecraft, and current desired location:
$\boldsymbol{s}_{i}=[\boldsymbol{x}^{H}_{m,i},\dot{\boldsymbol{x}}^{H}_{m,i},\boldsymbol{a}_{i}+\boldsymbol{x}^{H}_{m,i},{\boldsymbol{x}^{H}_{a,i-N}},...,{\boldsymbol{x}^{H}_{a,i}}]$ (6)
where the subscript $a$ represents the adversary or cat, and the subscript $m$ indicates the mouse. Although the derivative of the cat’s position is not explicitly provided, it is implicitly included in the time series data. The action $\boldsymbol{a}_{i}$ is then represented by the desired change in cartestian coordinates in the Hill frame:
$\boldsymbol{a}=[\Delta x^{H},\Delta y^{H},\Delta z^{H}]$ (7)
Additionally, the received state of the cat is a noisy observation $z:\mathcal{S}$ which is an estimate of the true location with gaussian noise:
${{\boldsymbol{z}_{a,i}}^{H}}=\mathcal{N}({\boldsymbol{x}^{H}_{a,i}},\Sigma_{t})$ (8)
Thus, the problem is a partially-observable markov decision process (POMDP). POMDPs are well studied and can be considerably more complex to learn an optimal policy for [NEURIPS2021_d5ff1353, DBLP:journals/corr/abs-2110-05038, DBLP:journals/corr/abs-2110-05038]. However, because the observed state $\boldsymbol{z}_{a,i}$ closely represents the true state, the system is assumed to behave as an MDP with gaussian noise in the observation – similar to many domain randomization methods [5979644, DBLP:journals/corr/abs-2110-03239, 8202133].
IV Simulation
In order to train a successful policy for evasion a simulation environment that accurately emulates the contested space environment as well as the sensor data collected is essential. In this section we define the framework for our simulation which models the orbital dynamics of each spacecraft, sensor constellation configurations, and finally RF localization.
IV-A Orbital Dynamics Modeling
The dynamic modeling of our simulation is composed of two layers: first the orbital dynamics of all satellites can be modeled using Kepler’s Laws of Planetary Motion, and second the localized dynamics of the cat and mouse spacecraft are modeled using the Clohessy Wiltshire Equations. All orbits are assumed to be near circular, and therefore each one can be defined by five parameters: inclination ($i$), argument of periapsis ($\omega$), right ascension of ascending node (RAAN) ($\Omega$), semi-major axis ($a$), and mean anomaly (M).
@@ IV-A Orbital Dynamics Modeling
Equations 9-14 allow us to solve for the position and velocity of the a spacecraft’s orbit at any point in time during the simulation.
In a localized frame the state of our cat and mouse spacecraft can be described using the Clohessy Wiltshire (CW) Equations. These equations model the behavior of a spacecraft in relation to a single point moving on a near circular orbit. We can track the location of this point using Equations 9-14. The equations of motion for the two spacecrafts position relative to this point is then:
@@
IV-B RF Modeling and Constellation Configuration
@@
$\Delta d_{i}=||\boldsymbol{x}^{E}_{a}-\boldsymbol{x}^{E}_{i}||-||\boldsymbol{x}^{E}_{a}-\boldsymbol{x}^{E}_{1}||$ (20)
In Equation 20 $\boldsymbol{x}^{E}_{1}$ is the location of the reference sensor.
IV-C Cat Localization
@@
Multiple works cited in Section II have shown their methods to perform similar to the CRLB both in simulation and real world testing [ZHOU2021109758, s23146254, 7582404, 10356280]. Thus, in order to reduce computational complexity of the simulation the CRLB is computed using Equations 21-24 in order to provide an appropriate noise level for the cat’s location estimate rather than run a TDOA localization algorithm using simulated TDOA data every timestep. The corresponding position estimate can be simulated by summing a sampled error using the CRLB and the true position of the cat:
${\boldsymbol{z}^{E}_{a}}=\boldsymbol{x}^{E}_{a}+\mathcal{N}(0,\,\sqrt{diag(\text{CRLB})})$ (25)
Due to the distance between the RF sensors and source as well as the fact that the sensors are constantly moving in relation to the source the size of the noise can vary from nominal to extreme. A graph depicting the error of the position estimates during a single episode can be seen in Figure 4. For this reason, we find it necessary to incorporate a filter into our system in order to properly process the observation data which will be discussed in Section V.
@@
V Method
In this section we will discuss our proposed control scheme of evasive maneuvering of the mouse. In order to properly evade an uncooperative actor while maintaining minimal deviation from the initial path an agent must be capable of perceiving its surroundings, processing observations, computing optimal actions, and finally executing said actions in a single feedback loop. Control of the spacecraft can be broken into two stages: the RL policy which receives an observation state of the environment and computes the a desired change in position, and an MPC controller (previously defined as $f(\boldsymbol{x}_{i},\dot{\boldsymbol{x}}_{i},\boldsymbol{a}_{i})$) which converts positional commands to thrust control. In addition, we implement an Extended Kalman Filter during testing in order to improve accuracy of the observations received and overall performance of the constrained RL policy.
V-A Reinforcement Learning Training Scheme
The RL policy is tasked with computing optimal positional control given some observation of the environment. The observation provided to the cat is a concatenation of the mouse’s state in the local frame, the mouse’s current action, and the estimated state of the cat in the local frame from the last $N$ timesteps:
$\boldsymbol{o}_{i}=[\boldsymbol{x}^{H}_{m,i},\dot{\boldsymbol{x}}^{H}_{m,i},\boldsymbol{a}_{i}+\boldsymbol{x}^{H}_{m,i},{\boldsymbol{z}^{H}_{a,i-N}},...,{\boldsymbol{z}^{H}_{a,i}}]$ (26)
As described by Equation 7 $\boldsymbol{a}_{i}$, is the desired change in current location of the mouse. This is then used to compute the desired location of the mouse in the local coordinate frame, and an MPC is utilized to compute the necessary thrust control to be applied.
In order to gauge the utility of an action during training, the policy is provided feedback in the form of a reward. During training, the reward function implemented prioritizes deviation from the initial orbit, fuel consumed, and proximity to the cat: